OpenStudy (anonymous):

Solve the following inequality and write your solution using set builder notation. 5/9(x) >= - 3/4 5/9(x) >= - 3/4 5/9-5/9(x) >= - 3/4 -5/9 x>= -3/4(9)-5/9(4) x>= -27/36 - 20/36 x>=-7/36

OpenStudy (anonymous):

{x|x>=7/36}

OpenStudy (a_clan):

\[(5/9) x \ge -3/4\] multiply 9/5 both sides, \[x \ge -27/20\] solution = { x | x ϵ R, x >= -27/20 }

OpenStudy (anonymous):

thanks. i turned it in .